OpenStudy (anonymous):

A world-class sprinter can burst out of the blocks to essentially top speed (of about 11.5m/s) in the first a5.0m of the race. What is the average acceleration of this sprinter, and how long does it take her to reach this speed?

OpenStudy (anonymous):

In general you need two standard equations: (1) Vt = Vo + a*t, with Vo the speed at start (here 0); (2) Xt = Xo + Vo*t + a/2*t^2 As Xo and Vo are 0, both equations reduce to: (1.1) Vt = a*t (1.2) Xt = a/2 * t^2 where Vt = 11.5 m/s and Xt = 5.0 m. This gives two equations with two unknowns (acceleration a and time-to-destination t). These are exactly the two variables asked for. Solve this system of equations and find a and t.